OpenStudy (anonymous):

7x-y=4 5x+2y=1 i know y=-13/19 but how do i plug that in to get x?

OpenStudy (anonymous):

u said that y=-13/19 so i am finding the value of x given : 7x-y=4 7x-(-13/19)=4 7x+13/10=4 7x=4-13/19 7x=63/19 19*7x=63 x=63/(19*7) x=9/19 VERIFY : 7x-y=4 7(9/19)-(-13/19)=4 63/19 +13/19 = 4 (63+13)/19=4 76/19=4 4=4 HENCE VERIFIED YOUR ANSWER IS X=9/19